# Distributed Switching Model Predictive Control Meets Koopman Operator for Dynamic Obstacle Avoidance

### v1

@@ 1 Introduction
Obstacle avoidance in dynamic environments demands both rapid response and reliable motion forecasting. Recent MPC-based methods have incorporated probabilistic and learning-based models for obstacle behavior (li2023moving; olcay2024dynamic; wei2022moving). Despite this progress, few efforts address online modeling of unknown dynamics for embedded control. This motivates the use of the Koopman operator, which linearly embeds nonlinear systems into higher-dimensional spaces, enabling efficient data-driven prediction (williams2015data). Koopman-based models have shown success in motion planning (gutow2020koopman), object tracking (comas2021self), robust control (zhang2022robust), and mobility applications (manzoor2023vehicular).
@@ 3 Main results
3.1 The proposed Koopman operator-based collision avoidance

### v2

@@ 1 Introduction
Obstacle avoidance in dynamic environments demands both rapid response and reliable motion forecasting. Recent MPC-based methods have incorporated probabilistic and learning-based models for obstacle behavior (li2023moving; olcay2024dynamic; wei2022moving). Despite this progress, few efforts address online modeling of unknown dynamics for embedded control. This motivates the use of the Koopman operator, which linearly embeds nonlinear systems into higher-dimensional spaces, enabling efficient data-driven prediction (williams2015data).
The Koopman operator has emerged as a powerful tool for bridging nonlinear dynamics and linear control frameworks. A comprehensive treatment of its theoretical underpinnings and relevance to systems and control is provided in (mauroy2020koopman), outlining its use in estimation, observability, and feedback design. The approach is utilized for controlled dynamical systems in (korda2018linear), where data-driven linear predictors are constructed in lifted spaces, enabling integration into efficient MPC schemes. The broader utility of the Koopman framework in forecasting, system identification, and controller synthesis is also surveyed in (mezic2021koopman). Koopman-based models have thus found success in motion planning (gutow2020koopman), object tracking (comas2021self), robust control (zhang2022robust), and mobility applications (manzoor2023vehicular).
@@ 3 Main results
3.1 Koopman operator-based collision avoidance